OpenStudy (anonymous):

Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard for, ax^2 +bx +c=0 the equation is _ =0

jimthompson5910 (jim_thompson5910):

You're missing the solutions.

OpenStudy (anonymous):

So sorry. 9 is the only solution

jimthompson5910 (jim_thompson5910):

x=9 or x=9 (note: if one solution is only present, then it's a double root) x-9=0 or x-9=0 (x-9)(x-9)=0 x^2-9x-9x+81=0 x^2-18x+81=0